# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ 3.2.3 Result for non-local operators
[…]
This explicitly ensures no spin-quantization occurs. The parameters $\alpha,\beta,\gamma$ and $\sigma$, which are implicitly $S$ and $\Delta$-dependent, will be part of the numerical output (see following Section 4).
We also noticed an extra constraint
${\alpha=\alpha^{\prime},\quad\alpha^{2}=\sec^{2}\left(\pi\Delta_{sl(2)}\right)}$ (149)
meaning that the independent parameters in each matrix are just three.
@@
Notice also that both these matrices reduce explicitly to the case for local operators if we demand that they lose the $u$-dependence, setting to zero the terms proportional to $e^{\pm\pi u}$, $e^{\pm 2\pi u}$. This can be achieved by setting, in the case of $\mathcal{G}_{\text{Regge}}$, $\gamma=\sigma=0$, and in the case of $\mathcal{G}_{\text{Bridge}}$, $\gamma^{\prime}=\sigma^{\prime}=0$. With these choices of parameters they both collapse to the form $\mathcal{G}_{\text{local}}$ (77).
Thus, the case of local operators sits precisely at the intersection of the cases described by the two more general gluing matrices.
@@
$\mathcal{G}_{\text{Bridge}}=-\mathbb{M}_{1}\cdot\mathcal{G}_{\text{Regge}}\cdot\mathbb{M}_{1}^{-1},$ (150)
upon an appropriate identification of the parameters $\alpha$, $\beta$, $\gamma$, $\sigma$, and $\alpha^{\prime}$, $\beta^{\prime}$, $\gamma^{\prime}$, $\sigma^{\prime}$ entering the gluing matrices on the two sides. Above, $\mathbb{M}_{1}$ is the matrix defined in (41), which realizes the Weyl symmetry $\texttt{W}_{S}$ at the level of the QSC. Notice that the minus sign on the r.h.s. of (150) has no observable effect, since the gluing matrix enters quadratically into the basic relation (17).
@@
$\mathcal{G}_{\text{Bridge}}=\mathbb{M}_{2}\cdot\mathcal{G}_{\text{Bridge}}\cdot\mathbb{M}_{2}^{-1},\;\;\;\;\mathcal{G}_{\text{Regge}}=\mathbb{M}_{2}\cdot\mathcal{G}_{\text{Regge}}\cdot\mathbb{M}_{2}^{-1}.$ (151)
Since $\mathbb{M}_{2}$ is a Q-system symmetry associated to the Weyl reflection in $\Delta$, this indicates that a trajectory with $\Delta$ and one with $\Delta\rightarrow d-\Delta$ are obtained by imposing the same form of gluing equations, just with redefined parameters.
As was anticipated in the Introduction, the existence of two distinct gluing matrices corresponds to two separate directions for performing the analytical continuation in spin, starting from physical local operators, and – by this very construction – the solution obtained with $\mathcal{G}_{\text{Bridge}}$ can be interpreted as a solution obtained with $\mathcal{G}_{\text{Regge}}$ subject to the Weyl transformation $\texttt{W}_{S}$. This will lead us straight to the existence of the twist/co-twist symmetry. The evidence for this conclusion is discussed in detail in the section 4.
3.3 Two gluing matrices also in $\mathcal{N}=4$ SYM!
@@
$\tilde{\mathbf{Q}}_{i}(u)=(\mathcal{G}^{\mathcal{N}=4})_{i}^{\,j}\cdot\bar{\mathbf{Q}}_{j}(u)\;\;\text{ ( SYM case ) }.$ (152)
@@
where $\xi=\xi_{1}+\xi_{2}\cosh(2\pi u)+\xi_{3}\sinh(2\pi u)$, with parameters $\chi,\xi_{1},\xi_{2}\in{\mathbb{C}}$ and $\nu\in\mathbb{R}$.
This form was used to compute the Regge trajectories with the same quantum numbers as the Pomeron in Gromov:2015wca ; Gromov:2017blm .
Notice that, just like in the ABJM case, when this matrix collapses to the constant case, i.e. when $\xi_{2}=\xi_{3}=0$, we find the form of the gluing conditions relevant for local operators (in this case one also has automatically $\gamma=\nu=0)$.
@@
$\mathcal{G}^{\mathcal{N}=4}_{\text{Bridge}}\simeq\left(\mathcal{G}^{\mathcal{N}=4}_{\text{Regge}}\right)^{T},$ (157)
@@ 4 Results
In this section we present our numerical studies. After a brief review of the numerical method which was first developed in Gromov:2015wca we present studies of the minimum of the leading trajectories interpolating odd and even spins. In Section 4.5 we present numerical evidence on the twist-co-twist symmetry. We describe in detail how it operates on trajectories and how it is related to the two gluing matrices found in the previous section.
@@ 4.1 Setup
$\Delta_{\text{top}}=\frac{\hat{M}_{1}+\hat{M}_{2}-3}{2},\;\;\;S_{\text{top}}=\frac{\hat{M}_{2}-\hat{M}_{1}-1}{2}.$ (162)
From this, the conformal charges (and R-symmetry) for any other representative of the multiplet, related to the top component by the action of supercharges, can be reconstructed by representation theory. In particular, we will often report results in the familiar $sl(2)$-like grading, which is related to the top component by the shifts (37).
@@ 4.2 Recap on the numerical method
The idea of the method follows the logical flow of equations (7), (9): the $\mathbf{P}$ functions are used as in input, parametrised in terms of some numerical constants– one then constructs $\mathbf{Q}$ functions using QQ relations and imposes that they should satisfy the correct gluing equation. Imposing the gluing equations gives a nonlinear condition which fixes the original parameters on a discrete set of solutions of the QSC.
@@
$\mathbf{P}_{ab}(u)=x^{\mathcal{N}_{a}+\mathcal{N}_{b}}\left(\mathcal{A}_{ab}+\sum_{n=1}^{\infty}\frac{c_{ab,n}}{x^{n}}\right),$ (163)
where $\left\{\mathcal{A}_{ab},c_{ab,n}\right\}$ are parameters and $x\equiv x(u)$ is the Zhukovsky map:
$x(u)=\frac{u+\sqrt{u-2h}\sqrt{u+2h}}{2h},$ (164)
which is a two-sheeted function with a single cut and asymptotics $x\sim u/h$ on the first sheet. The series (163) converges for $u$ on the first Riemann sheet and in a finite region of the second sheet. This depends on the fact that branch points of $\mathbf{P}$ functions are quadratic and thus are resolved by the Zhukovsky map. [^24] The lack of this property is what it makes it more difficult to solve the QSC for the AdS3 case: however see Cavaglia:2022xld for a way around this problem which works in some regimes.
@@
There is then a well-developed numerical procedure (see Gromov:2015wca ; Bombardelli:2018bqz for details) to construct the basis of functions $\mathbf{Q}^{\downarrow}_{ij}$ and $\mathbf{Q}^{\uparrow}_{ij}$, defined as solutions of the QQ relations (7) and (9). The numerical method proceeds by first computing the coefficients of the asymptotic expansion (18) or (19), which will depend on the input parameters; these asymptotic expansions then allow to evaluate the functions very far from the real axis in their respective half plane of holomorphicity; using the QQ relations, one can move towards the real axis in shifts of $-i$ or $+i$, and finally get to reconstructing the functions on the real axis. In particular, thanks to the Zhukovsky parametrisation we have easy access to these functions both above and below the branch cut $u\in[-2h,2h]$, see Gromov:2015wca ; Bombardelli:2018bqz for details.
@@
$\mathbf{Q}^{\downarrow}_{ij}(u\;;\;\texttt{params}),\;\;\;\mathbf{Q}^{\uparrow}_{ij}(u\;;\;\texttt{params}),\;\;\;\tilde{\mathbf{Q}}^{\downarrow}_{ij}(u\;;\;\texttt{params}),\;\;\;\tilde{\mathbf{Q}}^{\uparrow}_{ij}(u\;;\;\texttt{params}).$ (166)
@@
$\texttt{COST}\left[\texttt{params}\right]\equiv\sum_{k=1}^{N_{\text{points}}}\Big{|}\tilde{\mathbf{Q}}(u_{k};\texttt{params})-\left(\mathcal{G}\cdot\mathbf{Q}^{\uparrow}\cdot\mathcal{G}^{T}\right)(u_{k},\texttt{params})\Big{|}^{2}.$ (167)
@@
Typical values for the truncations used to generate the plots in this work are $N_{\text{trunc}}=22$ for the leading trajectory at $h=27/100$. We estimate the errors in the solution found in this way by truncating the last floating digit at the first iteration in which the error runs aground (see Appendix E for a sample of our results).

### v3

@@ 3.2.3 Result for non-local operators
    In particular, we find
    $g_{11}=g_{14}=g_{22}=g_{32}=g_{41}=g_{12}=g_{31}=g_{34}=g_{42}=0,$ (121)
    along with functional dependencies of the form $g_{ij}=g_{ij}(A,B,C,\hat{\mathcal{N}})$ for certain pairs $(i,j)$, as well as specific relations among the parameters $A$, $B$, $C$, and $\hat{\mathcal{N}}$.
  • Although the previous constraint (120) incorporates the quadratic condition (76), the latter is only used in the limit $u\rightarrow\pm\infty$. As a final step, we impose the constraint (76) on the exact form of $\mathcal{G}(u)$. The use of condition (120) for $u\rightarrow\pm\infty$ before imposing (76) on the exact gluing matrix is merely a convenience in the computation of the constraints. Alternative orderings of the algorithmic steps are possible, but they prove to be less advantageous for implementation. Finally we get
    $g_{22}^{(0)}=\mathrm{i},\quad g_{32}^{(0)}=0$ (122)
    together with some supplemental relations $g_{ij}=g_{ij}(A,B,C,\hat{\mathcal{N}})$ and between the coefficients $A,B,C$ and $\hat{\mathcal{N}}$. E.g. we find the constraint
    $\text{e}^{\mathrm{i}\pi\hat{\mathcal{N}}_{2}}=\mp\mathrm{i}\left(1+A_{2}^{+}B_{3}^{+}\right),$ (123)
    which replaces the condition (89): in the case of non-constant gluing matrix with $A_{2}^{+}$, $A_{3}^{+}\neq 0$, we see that the spin is no longer quantized. Furthermore, the following condition still holds,
    $(-1)^{M_{1}+M_{2}}=-1.$ (124)
Ultimately, we can parametrize the gluing matrix using only three parameters, $\beta,\gamma,\sigma\in\mathbb{C}$, [^22] The $\pm$ sign in this equation is correlated to the trajectory passing through even or odd integer spin local operators, in accordance with (88). Notice that, while an overall sign in $\mathcal{G}$ is invisible at the level of the gluing equations, definitions such as (79) pick a definite choice of sign.
±$\pm\left(\mathcal{G}_{\text{Regge}}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\alpha&0&0&-\frac{\mathrm{i}\left(\alpha^{2}-1\right)}{\beta}\\ 0&\mathrm{i}&\sigma&0\\ 0&0&-\mathrm{i}&0\\ \mathrm{i}\beta&0&0&\mathrm{i}\alpha\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&0&\gamma&0\\ -\frac{\beta\gamma\rho}{\alpha^{2}-1}&0&0&\gamma\frac{1-\rho}{\alpha}\\ 0&0&0&0\\ 0&0&-\frac{\beta\gamma\left(\alpha^{2}+\rho-1\right)}{\alpha\left(\alpha^{2}-1\right)}&0\\ \end{array}\right)+e^{-\pi u}$ (133)
×$\times\left(\begin{array}[]{cccc}0&0&\gamma\frac{\rho-1}{\alpha}&0\\ -\frac{\beta\gamma\left(\alpha^{2}+\rho-1\right)}{\alpha\left(\alpha^{2}-1\right)}&0&0&-\gamma\\ 0&0&0&0\\ 0&0&-\frac{\beta\gamma\rho}{\alpha^{2}-1}&0\\ \end{array}\right)+\left(e^{2\pi u}+e^{-2\pi u}\right)\times\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&-\frac{\mathrm{i}\beta\gamma^{2}(\rho-1)}{\alpha\left(\alpha^{2}-1\right)}&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{array}\right),$ (142)
where $\rho^{2}=1-\alpha^{2}\equiv 1-\text{csc}^{2}\left(\pi\hat{\mathcal{N}_{1}}\right)$.
If we instead use the other assumption, the “Bridge” case (94), the gluing matrix emerging from the algorithm can be formally taken to be the transpose of the above one, with its own set of parameters, $\beta^{\prime},\gamma^{\prime},\sigma^{\prime}\in\mathbb{C}$,
±$\pm\left(\mathcal{G}_{\text{Bridge}}\right)_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\alpha^{\prime}&0&0&\mathrm{i}\beta^{\prime}\\ 0&\mathrm{i}&0&0\\ 0&\sigma^{\prime}&-\mathrm{i}&0\\ -\frac{\mathrm{i}\left(\alpha^{\prime 2}-1\right)}{\beta^{\prime}}&0&0&\mathrm{i}\alpha^{\prime}\\ \end{array}\right)+e^{\pi u}\times\left(\begin{array}[]{cccc}0&-\frac{\beta^{\prime}\gamma^{\prime}\rho^{\prime}}{\alpha^{\prime 2}-1}&0&0\\ 0&0&0&0\\ \gamma^{\prime}&0&0&-\frac{\beta^{\prime}\gamma^{\prime}\left(\alpha^{\prime 2}+\rho^{\prime}-1\right)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}\\ 0&\gamma^{\prime}\frac{1-\rho^{\prime}}{\alpha^{\prime}}&0&0\\ \end{array}\right)+e^{-\pi u}$ (151)
×$\times\left(\begin{array}[]{cccc}0&-\frac{\beta^{\prime}\gamma^{\prime}\left(\alpha^{\prime 2}+\rho^{\prime}-1\right)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}&0&0\\ 0&0&0&0\\ \gamma^{\prime}\frac{\rho^{\prime}-1}{\alpha^{\prime}}&0&0&-\frac{\beta^{\prime}\gamma^{\prime}\rho^{\prime}}{\alpha^{\prime 2}-1}\\ 0&-\gamma^{\prime}&0&0\\ \end{array}\right)+\left(e^{2\pi u}+e^{-2\pi u}\right)\times\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&-\frac{\mathrm{i}\beta^{\prime}\gamma^{\prime 2}(\rho^{\prime}-1)}{\alpha^{\prime}\left(\alpha^{\prime 2}-1\right)}&0&0\\ 0&0&0&0\\ \end{array}\right),$ (160)
where $\left(\rho^{\prime}\right)^{2}=1-\left(\alpha^{\prime}\right)^{2}\equiv 1-\text{csc}^{2}\left(\pi\hat{\mathcal{N}_{1}}\right)$. The parameters $\beta,\gamma$ and $\sigma$, which are implicitly $S$ and $\Delta$-dependent, will be part of the numerical output (see following Section 4).
@@
Notice also that both these matrices reduce explicitly to the case for local operators if we demand that they lose the $u$-dependence, setting to zero the terms proportional to $e^{\pm\pi u}$, $e^{\pm 2\pi u}$. This can be achieved by setting, in the case of $\mathcal{G}_{\text{Regge}}$, $\gamma=\sigma=0$, and in the case of $\mathcal{G}_{\text{Bridge}}$, $\gamma^{\prime}=\sigma^{\prime}=0$. With these choices of parameters they both collapse to the form $\mathcal{G}_{\text{local}}$ (88). Thus, the case of local operators sits precisely at the intersection of the cases described by the two more general gluing matrices.
@@
$\mathcal{G}_{\text{Bridge}}\cong-\mathbb{M}_{1}\cdot\mathcal{G}_{\text{Regge}}\cdot\mathbb{M}_{1}^{-1},$ (161)
upon an appropriate identification of the parameters $\alpha$, $\beta$, $\gamma$, $\delta$ and $\alpha^{\prime}$, $\beta^{\prime}$, $\gamma^{\prime}$, $\delta^{\prime}$ entering the gluing matrices on the two sides. Above, $\mathbb{M}_{1}$ is the matrix defined in (49), which realizes the Weyl symmetry $\texttt{W}_{S}$ at the level of the QSC. Notice that the minus sign on the r.h.s. of (161) has no observable effect, since the gluing matrix enters quadratically into the basic relation (17).
As was anticipated in the Introduction, the existence of two distinct gluing matrices $\mathcal{G}_{\text{Regge}}$ and $\mathcal{G}_{\text{Bridge}}$ corresponds to two separate directions for performing the analytical continuation in spin starting from physical local operators. Since the two are related by the transformation (161), any solution obtained with $\mathcal{G}_{\text{Bridge}}$ can be interpreted as a solution obtained with $\mathcal{G}_{\text{Regge}}$ subject to the Weyl transformation $\texttt{W}_{S}$. This will lead us straight to the existence of the twist/co-twist symmetry. The evidence for this conclusion is discussed in detail in the Section 4.
@@
$\mathcal{G}_{\text{Regge}}\cong\mathbb{M}_{2}\cdot\mathcal{G}_{\text{Regge}}\cdot\mathbb{M}_{2}^{-1},\quad\mathcal{G}_{\text{Bridge}}\cong\mathbb{M}_{2}\cdot\mathcal{G}_{\text{Bridge}}\cdot\mathbb{M}_{2}^{-1}.$ (162)
Since $\mathbb{M}_{2}$ is a Q-system symmetry associated to the Weyl reflection in $\Delta$, this indicates that a solution with $\Delta$ and one with $\Delta\rightarrow d-\Delta$ are obtained by imposing the same form of gluing equations, just with redefined parameters.
3.3 Two gluing matrices also in $\mathcal{N}$$=$$4$ SYM
@@
[…]
@@
where $\xi=\xi_{1}+\xi_{2}\cosh(2\pi u)+\xi_{3}\sinh(2\pi u)$, with parameters $\chi,\xi_{1},\xi_{2}\in{\mathbb{C}}$ and $\nu\in\mathbb{R}$. This form was used to compute the Regge trajectories with the same quantum numbers as the Pomeron in Gromov:2015wca ; Gromov:2017blm . Notice that, just like in the ABJM case, when this matrix collapses to the constant case, i.e. when $\xi_{2}=\xi_{3}=0$, we find the form of the gluing conditions relevant for local operators (in this case one also has automatically $\gamma=\nu=0)$.
@@
$\mathcal{G}^{\mathcal{N}=4}_{\text{Bridge}}\cong\left(\mathcal{G}^{\mathcal{N}=4}_{\text{Regge}}\right)^{T},$ (168)
@@
Beyond the LR symmetric sector, in $\mathcal{N}$$=$$4$ SYM there exists a second Weyl reflection corresponding to $S_{2}\rightarrow-S_{2}$. This too should correspond to a further transformation for the gluing matrix. Alternative gluing matrices related to $S_{2}$ and $-S_{2}$ were indeed found by the authors of Alfimov:2018cms , where they were used to move in a neighbourhood of the point $S_{2}=0$ in two directions. The analogue of what we are discussing in this paper would be to use the alternative gluing matrix around a generic physical operator with $S_{2}\neq 0$, to generate a new $S_{2}$-flipped trajectory intersecting the first one, thus highlighting a new symmetry. [^24] We thank Nikolay Gromov for communications.
@@ 4 Results
In this section we present our numerical studies. After a brief review of the numerical method which was first developed in Gromov:2015wca we present studies of the minimum of the leading trajectories interpolating odd and even spins. In Section 4.5 we present numerical evidence on the twist/co-twist symmetry. We describe in detail how it operates on trajectories and how it is related to the two gluing matrices found in the previous section.
@@ 4.1 Setup
$\Delta_{\text{top}}=\frac{\hat{M}_{1}+\hat{M}_{2}-3}{2},\quad S_{\text{top}}=\frac{\hat{M}_{2}-\hat{M}_{1}-1}{2}.$ (173)
From this, the conformal charges (and R-symmetry) for any other representative of the multiplet can be reconstructed by acting with the appropriate supercharges. We will often report results in the familiar $sl(2)$-like grading, which is related to the charges of the top component by the shifts (37).
@@ 4.2 Recap on the numerical method
The idea of the method follows the logical flow of equations (7), (9): the $\mathbf{P}$ functions are used as in input, parametrised in terms of some numerical constants– one then constructs $\mathbf{Q}$ functions using QQ-relations and imposes that they should satisfy the correct gluing equation. Imposing the gluing equations gives a nonlinear condition which fixes the original parameters on a discrete set of solutions of the QSC.
@@
$\mathbf{P}_{ab}(u)=x^{\mathcal{N}_{a}+\mathcal{N}_{b}}\left(\mathcal{A}_{ab}+\sum_{n\geq 1}\frac{c_{ab,n}}{x^{n}}\right),$ (174)
where $\left\{\mathcal{A}_{ab},c_{ab,n}\right\}$ are parameters and $x\equiv x(u)$ is the Zhukovsky map (66), which we repeat here for clarity:
$x(u)=\frac{u+\sqrt{u-2h}\sqrt{u+2h}}{2h}.$ (175)
This map is a two-sheeted function with a single cut and asymptotics $x\simeq u/h$ on the first sheet. The series (174) converges for $u$ on the first Riemann sheet and in a finite region of the second sheet. This depends on the fact that branch points of $\mathbf{P}$ functions are quadratic and thus are resolved by the Zhukovsky map. [^27] The lack of this property is what it makes it more difficult to solve the QSC for the AdS3 case: however see Cavaglia:2022xld for a way around this problem which works in some regimes.
@@
There is then a well-developed numerical procedure (see Gromov:2015wca ; Bombardelli:2018bqz for details) to construct the basis of functions $\mathbf{Q}^{\downarrow}_{ij}$ and $\mathbf{Q}^{\uparrow}_{ij}$, defined as solutions of the QQ-relations (7) and (9). The numerical method proceeds by first computing the coefficients of the asymptotic expansion (18) or (19), which will depend on the input parameters; these asymptotic expansions then allow to evaluate the functions very far from the real axis in their respective half plane of holomorphicity; using the QQ-relations, one can move towards the real axis in shifts of $-\mathrm{i}$ or $\mathrm{i}$, and finally get to reconstructing the functions on the real axis. In particular, thanks to the Zhukovsky parametrisation we have easy access to these functions both above and below the branch cut $u\in[-2h,2h]$, see Gromov:2015wca ; Bombardelli:2018bqz for details.
@@
$\mathbf{Q}^{\downarrow}_{ij}(u\,;\,\texttt{params}),\quad\mathbf{Q}^{\uparrow}_{ij}(u\,;\,\texttt{params}),\quad\tilde{\mathbf{Q}}^{\downarrow}_{ij}(u\,;\,\texttt{params}),\quad\tilde{\mathbf{Q}}^{\uparrow}_{ij}(u\,;\,\texttt{params}).$ (177)
@@
$\texttt{COST}\left[\texttt{params}\right]\equiv\sum_{k=1}^{N_{\text{points}}}\Big{|}\tilde{\mathbf{Q}}^{\downarrow}(u_{k};\texttt{params})-\left(\mathcal{G}\cdot\mathbf{Q}^{\uparrow}\cdot\mathcal{G}^{T}\right)(u_{k};\texttt{params})\Big{|}^{2}.$ (178)
@@
For our numerical study, we typically used a truncation parameter $N_{\text{trunc}}\simeq 20$, which allows us to reach a precision of e.g. $\simeq 20-30$ digits for the leading trajectory at coupling $h=0.27$ (some precision is lost around the pomeron point). The primary source of error in the solution obtained through this approach arises from the systematic bias introduced by truncations. We estimate the error by discarding the digits affected by increasing the truncation parameters. See Appendix E for a representative sample of our results.